OpenStudy (anonymous):

The temperature is changing every day by negative 3 over 16 degree Fahrenheit. What will be the change in the temperature after 4 days? negative 64 over 3 degrees Fahrenheit negative 3 over 4 degree Fahrenheit negative 3 over 64 degree Fahrenheit 3 over 4 degree Fahrenheit

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

(TEMPERATURE CHANGE) The rate at which the temperature of an object changes is proportional to the difference between its own temperature and the temperature of the surrounding medium. Express this rate as a function of the temperature of the object

OpenStudy (anonymous):

Hint

OpenStudy (anonymous):

I think it is C

OpenStudy (skullpatrol):

Why C?

OpenStudy (anonymous):

cuz 4 X 16 = 64

OpenStudy (anonymous):

lol

OpenStudy (skullpatrol):

$$\huge-3/16 + (-3/16) + (-3/16) + (-3/16) = 4• (-3/16)$$

OpenStudy (anonymous):

O.O

OpenStudy (skullpatrol):

The change is -3/16 per day for 4 days.

OpenStudy (anonymous):

so wat is it?

OpenStudy (anonymous):

@TheCuteOne25

OpenStudy (skullpatrol):

4•(-3/16) = ?

OpenStudy (anonymous):

-0.75?

OpenStudy (skullpatrol):

yes

OpenStudy (thecuteone25):

yeah I got -0.75

OpenStudy (skullpatrol):

-0.75 = 75/100 = 3/4